OpenStudy (anonymous):

y = -6.9x + 5.1 A. -69x + 10y = 51 B. 69x + 10y = 51 C. 69x + 10y = -51 D. -69x + 10y = -51

OpenStudy (anonymous):

What is the question? what are you trying to do with the first equation?

OpenStudy (anonymous):

@Catseyeglint911

OpenStudy (anonymous):

What is the equation of the given line in standard form? Use integer coefficients.

OpenStudy (anonymous):

\[y=-6.9x + 5.1\] add 6.9 to each side: \[6.9x + y=5.1\] and since standard form does not include decimals multiply by 10 to get rid of them: \[69x+10y=51x\] and that should be your answer! :)

OpenStudy (anonymous):

First you need to add -6.9 to both sides 6.9x + y = 5.1 Next multiply both sides by 10 to get rid of the decimals. 69x +10y = 51 so B. 69x + 10y = 51

OpenStudy (anonymous):

ok thank you !!!!1